OpenStudy (anonymous):

Solve the equation, if possible. Give your answer in radical form. t^2=-81

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ a^2-b^2 = (a-b)(a+b)}} \\ \quad \\ \quad \\ t^2=-81\implies t^2+81=0\implies t^2+9^2=0\\ \quad \\ {\color{blue}{ (t-9)(t+9)}}=0\)

OpenStudy (jdoe0001):

ahemm rats... I got my signs mixed up... heeh. that's .. off.. one sec

OpenStudy (jdoe0001):

well, actually you'd get a complex anyhow if you just take root to both sides

OpenStudy (jdoe0001):

\(\bf t^2=-81\implies \sqrt{t^2}=\sqrt{-81}\implies t=\sqrt{-9^2}\\ \quad \\ t=\sqrt{9^2\cdot -1}\implies t=\sqrt{9^2}\cdot \sqrt{-1}\)

OpenStudy (anonymous):

is that in radical form??????

OpenStudy (triciaal):

-81 = -1 x 81 rt -81 = rt-1 rt 81 rt(t^2) = rt(-81) =(rt-1)(rt 81) rt-1 =i rt 81 = + or - 9

OpenStudy (anonymous):

Thank you to both of you!!!

OpenStudy (triciaal):

welcome

OpenStudy (jdoe0001):

yw

OpenStudy (mertsj):

\[t^2=-81\] \[t=\pm \sqrt{-81}\] \[t=\pm9i\]